The pressure of an ideal gas is reduced by 50%, resulting in a decrease in temperature to 75% of the initial value. Calculate the ratio of the final to initial volumes of the gas




1
Expert's answer
2021-06-21T11:40:17-0400
"\frac{V_f}{V_i}=\frac{p_i}{p_f}\frac{T_f}{T_i}\\
\frac{V_f}{V_i}=\frac{1}{0.5}0.75=1.5"